OpenStudy (anonymous):

Help: Find the inverse of the function. Is the inverse a function? y = (3x-4)^2 and y = (1-2x)^2 + 5

OpenStudy (anonymous):

solve for x : sqrt(y) = 3x-4 sqrt(y) + 4 = 3x x = (sqrt(y) + 4) /3 so the inverse : y = (sqrt(x) + 4) / 3

OpenStudy (anonymous):

just make sure : y = (+/-sqrt(x) + 4) / 3 so it is not a function!

OpenStudy (turingtest):

the other one: y-5=(1-2x)^2 sqrt(y-5)=1-2x 2x=1-sqrt(y-5) x=(1/2)(1-sqrt(y-5)) so to write inverse of the second function switch x for y: y=(1/2)(1-sqrt(x-5))

OpenStudy (anonymous):

Thank you very much!:)